18=9-(3-x)
show work

Found 2 solutions by tran3209, zeynep:
Answer by tran3209(100)
Answer by zeynep(43)

18=9-(3-x) Let's remove the parantheses.We will change the signs of all of the terms being subtracted.
18=9-3+x
18=6+x (Let's add additive inverse of 6 to both sides)
18-6=6+x-6
12=x
I hope it helps. If there is something you cannot understand do not hesitate to ask!